Wakefield Industries purchased a delivery vehicle on January 1, 2019, for $30,000. They paid sales taxes of $2,000 and one year’s insurance premium of $2,000 on the same day. The company estimated a residual value of $3,000 at the end of the 5-year expected service life.

Expected usage (in miles) of the new delivery vehicle
2019 16,000
2020 22,000
2021 25,000
2022 21,000
2023 16,000

Mileage estimates were based on expected increases in customers and the addition of a second delivery vehicle in 2022.

What was the book value of the vehicle at December 31, 2021, under the straight-line and units-of-production depreciation methods, respectively?

Select one:

a. $13,600; $12,580

b. $15,400: $14,470

c. $13,800; $12,990

d. $12,800; $11,840

e. $14,600; $13,730
Answer #1

Answer: e.$14,600; $13,730

Calculations:

Purchase of the vehicle $30,000
Sales tax $2,000
Total cost $32,000
(Less): Salvage value ($3,000)
Depreciable value $29,000

Straight-line Method;

  • Depreciation expense = Depreciable value ÷ Estimated useful life = $29,000/5 years = $5,800
  • Accumulated depreciation for 3 years [Jan 1,2019 to Dec 31,2021] = $5,800 x 3 years = $17,400
  • Book value at Dec 31, 2021 = Cost - Accumulated depreciation = $32,000 - $17,400 = $14,600

Units-of-production:

  • Depreciation Expense per unit = Depreciable value ÷ Total estimated units of production = $29,000/100,000 = $0.29
  • Accumulated depreciation for 3 years [Jan 1,2019 to Dec 31,2021] = [16,000 + 22,000 + 25,000] x $0.29 = 63,000 units x $ 0.29 = $18,270
  • Book value at Dec 31,2021 = Cost - Accumulated depreciation = $32,000 - $18,270 = $13,730

Thus, Option e is correct and remaining options are incorrect.